Circular No. 9025

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 N3 (LULIN) M. Knight and D. Schleicher, Lowell Observatory, obtained and an-

alyzed extensive CN narrowband images of comet C/2007 N3 on eleven nights between Jan. 30 and Mar. 2, along with CN snapshots on five additional nights, using the Hall 1.1-m telescope at Lowell Observatory. Following the removal of median radial profiles, two side-on gas jets are seen centered at position angles near 110^{o} and 290^{o}, with each showing a corkscrew morphology. Intervals between fifteen pairs of matching CN images from various rotational cycles, coupled with the spacing and the outward motion of the features, imply a nucleus rotation period of 42.0 ± 0.5 hr. Preliminary modeling yields an obliquity of the rotation axis near 95^{o} and source locations within 30^{o} of each pole.

COMETS C/2008 W1 AND C/2009 A1 (STEREO) Further to IAUCs 8955 and 9005, K. Battams has measured (and

B. G. Marsden has reduced) astrometry for two additional near-sun presumed comets that were found on STEREO website images by A. Watson. C/2008 W1, which was observed only by the ‘HI1-B’ instrument (onboard STEREO-B), was a Kreutz sungrazer. C/2008 W1 peaked at magnitude perhaps 9–10; the SOHO LASCO instrument was not operational when this object was found. C/2009 A1, which was detected by the HI1 instruments on both STEREO-A and STEREO-B, belongs to no known group; D. Chestnov and R. Kracht aided in the finding of images of this object in HI1-B images after it was initially found in HI1-A images. Battams notes that C/2009 A1 was very faint, peaking around mag 13 (with 14 being the limiting magnitude of the HI-1 instruments), and it was small and seemingly diffuse. The parabolic orbital elements below for C/2009 A1 by Marsden were published on MPC 65057.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 W1 Nov. 18.739 12^{h}28.^{m}7 − 7^{o}38′ H1B AW 2009-B10
Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 A1 Jan. 10.017 22^{h}28.^{m}1 − 3^{o}51′ H1A AW 2009-B10
T = 2009 Jan. 11.785 TT  ω = 102.^{o}326
Ω = 340.014 }2000.0
q = 0.12713 AU  i = 51.472
2009 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green